OpenStudy (lanadelbey):

Find the 6th term of the sequence with t1 = -4 and tn = 5tn-1.

OpenStudy (lanadelbey):

@jim_thompson5910 can you help me with this one?

jimthompson5910 (jim_thompson5910):

"t1 = -4" means \(\Large t_1 = -4\) which is a fancy way of saying "the first term is -4" agreed?

OpenStudy (lanadelbey):

agreed

jimthompson5910 (jim_thompson5910):

\[\Large t_n = 5*t_{n-1}\] is a mathematical way of saying nth term = 5*(term just before the nth term)

jimthompson5910 (jim_thompson5910):

in other words, it's instructions on how to generate each term. It's very similar to the last problem (instead of 5, replace it with -2)

jimthompson5910 (jim_thompson5910):

if -4 is the first term, what is the second term?

OpenStudy (lanadelbey):

so it'd be -4*5 = -20?

jimthompson5910 (jim_thompson5910):

correct, -20 is the second term

OpenStudy (lanadelbey):

so the 6th term is -12,500

jimthompson5910 (jim_thompson5910):

good